OpenStudy (anonymous):

a radioactive isotope has a half-life of 2 weeks. if the original of the radiation is 1000 counts per second, how many weeks must elapse until the strength ot the radiation is less than 1 count per second

OpenStudy (anonymous):

1000*(1/2)^x = 1, where x is how many 'cycles' of 2 weeks you need to wait. I get x = 9.965, so you'd have to wait about 19 weeks

OpenStudy (anonymous):

how do you get the 19 weeks

OpenStudy (anonymous):

you do the value of x, so 9.965 * 2 = 19.93. should've said about 20 week, my mistake, but that is how you get the number of weeks

OpenStudy (anonymous):

thanks. so the equation is the one at the top

OpenStudy (anonymous):

please can you come back here

OpenStudy (anonymous):

what did you want to know?

OpenStudy (anonymous):

how did you get the 9.965

OpenStudy (anonymous):

i solved for x in the equation 1000*(1/2)^x = 1. i had to use a calculator, but you can also use logarithms.

OpenStudy (anonymous):

to do that you would do: (1/2)^x = 1/1000 ln ((1/2)^x) = ln(1/1000) x*ln ((1/2)) = ln(1/1000) x = \(ln (1/2)/ln(1/1000)\)

OpenStudy (anonymous):

oh sorry, i should have said: x = ln(1/1000)/ln (1/2)

OpenStudy (anonymous):

where 'ln' is the natural logarithm

OpenStudy (anonymous):

than I multiply 9.9657by 2 is that right

OpenStudy (anonymous):

yes, that will give you how long you would have to wait

OpenStudy (anonymous):

to get a count rate of 1

OpenStudy (anonymous):

ok understood.